# Your Dense Retriever is Secretly an Expeditious Reasoner

### v1

@@ 1 Introduction
Nevertheless, certain queries may not be adequately addressed by the structured transformation of dense reasoning. In such cases, retaining LLM-based reasoning is essential to ensure robust overall retrieval performance, thus a routing mechanism is needed to ensure appropriate reasoning process (Bai et al., 2024; Zhang et al., 2025a). To this end, we introduce the Reasoner Router ($R$). Given a new query, the Reasoner Router determines whether it can be reliably handled by the Dense Reasoner, i.e., whether the learned structured transformation can stably approximate the semantic effect of LLM reasoning, and otherwise resorts to LLM for deep reasoning. This mechanism enables a controllable trade-off between efficiency and accuracy. Built on the components described above, we propose the Adaptive Query Reasoning (AdaQR) framework. AdaQR is a hybrid reasoning pipeline consisting of three components: an LLM Reasoner, a Dense Reasoner, and a Reasoner Router. The Reasoner Router flexibly directs each user query to either the Dense Reasoner or the LLM Reasoner, preserving—and in some cases improving—the retrieval quality achieved by full LLM rewrites, while substantially reducing the reasoning (see Figure 1, right).
@@ 3.2 Main Result
Different dense retrievers have a significant impact on AdaQR. For ReasonIR-8B, AdaQR delivers the most outstanding and stable performance improvement, achieving an average cost reduction of approximately 29.21% while improving nDCG performance by about 13.18%. In contrast, Qwen3-Embedding-0.6B and Qwen3-Embedding-4B show poor performance. When using these models, AdaQR delivers very limited performance gains and even exhibits slight performance decrements in some few scenarios. This phenomenon aligns with the MRL results in Section $2.$ Embedding transformation for dense retrievers with high MRL values, such as ReasonIR-8B, can be more effectively learned, thus significantly enhancing AdaQR’s performance.
@@ 3.4 Ablation Study
To explore the individual contributions of the Dense Reasoner and the Reasoner Router, two core components within our proposed AdaQR framework, we conduct a comprehensive ablation study. Specifically, we remove Reasoner Router component, and further removed both Dense Reasoner and Reasoner Router (directly using the original query without any reasoning). Figure 5 reports the results of our ablation study: the average nDCG of AdaQR, w/o $R$ and w/o DR+$R$ are 21.05, 17.37 and 15.59 respectively, which clearly reveal the necessity of both components.
@@ 3.5 Effect of Decision Threshold $\tau$
It is worth noting that the strategy represented by w/o $R$ offers a near-zero-cost query reasoning method while still outperforming the original query approach, offering a promising option in scenarios with extremely constrained computational resources.

### v2

@@ 1 Introduction
Nevertheless, certain queries may not be adequately addressed by the structured transformation of dense reasoning. In such cases, retaining LLM-based reasoning is essential to ensure robust overall retrieval performance, thus a routing mechanism is needed to ensure appropriate reasoning process (Bai et al., 2024; Zhang et al., 2025a). To this end, we introduce the Reasoner Router (RR). Given a new query, the Reasoner Router determines whether it can be reliably handled by the Dense Reasoner, i.e., whether the learned structured transformation can stably approximate the semantic effect of LLM reasoning, and otherwise resorts to LLM for deep reasoning. This mechanism enables a controllable trade-off between efficiency and accuracy. Built on the components described above, we propose the Adaptive Query Reasoning (AdaQR) framework. AdaQR is a hybrid reasoning pipeline consisting of three components: an LLM Reasoner, a Dense Reasoner, and a Reasoner Router. The Reasoner Router flexibly directs each user query to either the Dense Reasoner or the LLM Reasoner, preserving—and in some cases improving—the retrieval quality achieved by full LLM rewrites, while substantially reducing the reasoning (see Figure 1, right).
@@ 3.2 Main Result
Different dense retrievers have a significant impact on AdaQR. For ReasonIR-8B, AdaQR delivers the most outstanding and stable performance improvement, achieving an average cost reduction of approximately 29.21% while improving nDCG performance by about 13.18%. In contrast, Qwen3-Embedding-0.6B and Qwen3-Embedding-4B show poor performance. When using these models, AdaQR delivers very limited performance gains and even exhibits slight performance decrements in some few scenarios. This phenomenon aligns with the MRL results in Section 2.2. Embedding transformation for dense retrievers with high MRL values, such as ReasonIR-8B, can be more effectively learned, thus significantly enhancing AdaQR’s performance.
@@ 3.4 Ablation Study
To explore the individual contributions of the Dense Reasoner and the Reasoner Router, two core components within our proposed AdaQR framework, we conduct a comprehensive ablation study. Specifically, we remove Reasoner Router component, and further removed both Dense Reasoner and Reasoner Router (directly using the original query without any reasoning). Figure 5 reports the results of our ablation study: the average nDCG of AdaQR, w/o RR and w/o DR+RR are 21.05, 17.37 and 15.59 respectively, which clearly reveal the necessity of both components.
@@ 3.5 Effect of Decision Threshold $\tau$
It is worth noting that the strategy represented by w/o RR offers a near-zero-cost query reasoning method while still outperforming the original query approach, offering a promising option in scenarios with extremely constrained computational resources.